OpenStudy (anonymous):

arctan(2x)+tan(x)=pi/4, what is x?

hartnn (hartnn):

arctan(2x)+tan(x)=pi/4 ?? or arctan(2x)+arctan(x)=pi/4

hartnn (hartnn):

if it is latter, then use arctanA+arctanB = arctan((A+B)/(1-AB)) then take tan on both sides. tan pi/4=1

hartnn (hartnn):

u get 3x=1-2x^2 solve this quadatic

OpenStudy (anonymous):

thanks!! i had solve it now :)) could you answer the other question i had closed?? the one with the round holes?